OpenStudy (anonymous):

if 20 men or 24 women or 40 boys can do a job in 12 days working for 8hours a day.how many men working with 6 women and 2 boys take to do a job four times as big working 5 hours a day for 12 days?

OpenStudy (anonymous):

26?

OpenStudy (mathstudent55):

To do the original job, it takes: Men: 20 men * 96 hours = 1920 man-hours, or Women: 24 women * 96 hours = 2304 woman-hours, or Boys: 40 boys * 96 hours = 3840 boy-hours The new job is 4 times as big, so it will take Men: 7680 man-hours, or Women: 9216 woman-hours, or Boys: 15,360 boy-hours In 5 days of 12-hour days there are 60 hours of work. 6 women do 6 * 60 hours = 360 woman-hours, and 2 boys do 2 * 60 hours = 120 boy-hours of work For the new job: The women did 360/9216 = 0.0390625 part of the new job The boys did 120/15,360 = 0.0078125 part of the new job The work left to do is 1 - 0.0390625 - 0.0078125 = 0.953125 part of the job left for the men to do. Since the men do the entire new job in 7680 man-hours, 0.953125 part of the job will take 0.953125 * 7680 = 7320 man-hours to do. Since the men can work for 60 hours, 7320/60 = 122. The answer is 122 men.

OpenStudy (anonymous):

oops didn't read the "4 times as big part" 26 was the answer if it's the same job, 1/(20*8*12) * n * 5* 12+ 1/(24*8*12) * (6) * 5 * 12 + 1/(40*8*12) * 2*5 * 12 = 4 n = (indeed) 122

OpenStudy (anonymous):

given answer is 2mens....

OpenStudy (anonymous):

I dont know how it is....

OpenStudy (mathstudent55):

2 men?

OpenStudy (anonymous):

that's just impossible even without doing any calculation

OpenStudy (zarkon):

maybe they meant 2 Zarkons. Two Zarkons could do the work, but not two regular men. If you just have regular men then you need 122 of them.

OpenStudy (anonymous):

LOL

OpenStudy (zarkon):

I'm just stating facts...no jokes here

OpenStudy (zarkon):

not sure how many sourwing's it would take...I'll have to think about that

OpenStudy (anonymous):

it would take infinite lol

OpenStudy (zarkon):

I'm betting somewhere between 2 and 122

OpenStudy (zarkon):

you are too modest